# revtex4-2Repair the float

### v1

@@ Abstract
Here we propose another approach that, similarly to the Ginzburg-Landau (GL) theory, relies solely on symmetry arguments. This allows us to incorporate effects of spin-orbit coupling into the NLSM and formulate a phenomenological quantum kinetic theory. Our theory captures all symmetry-allowed transport effects associated with spin, charge, and their coupling, such as spin-Hall and spin-galvanic effects. In contrast to GL, this phenomenology is valid for arbitrary temperatures, away from equilibrium, and covers the normal state. In the following, we identify the basic symmetries and explain how to construct the generalized NLSM. Then, the saddle point condition leads us to our main objective: obtaining the general quantum kinetic equation for disordered conductors with spin-orbit coupling, which reveals all physical effects associated with the different terms of the action.
@@
-- $\varepsilon_{jkl}\frac{D}{4}\theta\sigma_{i}Q\partial_{j}Q\partial_{k}Q+i\frac{D}{4}\varepsilon_{jkl}\varkappa\partial_{j}Q\partial_{k}Q+\frac{\Gamma}{8}\sigma_{k}Q\sigma_{k}Q\Big{)},$ (1)
where the Pauli matrices $\sigma_{i}$ and $\tau_{i}$ span the spin and the Nambu spaces, respectively, $\hat{\omega}_{tt^{\prime}}=\partial_{t}\delta(t-t^{\prime})$ is the time derivative operator, $D$ is the diffusion coefficient, and $\hat{\Delta}=\tau_{2}|\Delta|e^{i\tau_{3}\varphi}$ is the superconducting order parameter matrix. The second line in Eq. (1) describes the effects of extrinsic SOC: the spin Hall effect, spin current swapping and the spin relaxation, parameterized by the Hall angle $\theta$, the swapping coefficient $\varkappa$, and the spin relaxation rate $\Gamma$, respectively. The external electromagnetic potential is introduced through the replacement $\partial_{\mu}\mapsto\partial_{\mu}-i[A_{\mu}\tau_{3},.]$, and the Zeeman or exchange field is denoted by $\bm{h}$.
By inspecting the structure of $S[Q]$ in Eq. (1), we can observe the following. In general, the action represents a low-energy gradient expansion combined with an expansion in powers of spin Pauli matrices $\sigma_{i}$, and involves $\tau_{3}$ in the terms/fields breaking the time-reversal symmetry. Each term in $S$ is a scalar composed of $Q$, $\partial_{k}Q$, and $\sigma_{i}$, and respects the spatial symmetry of the system. In fact, the terms present in Eq. (1) exhaust all scalar rotation invariant forms up to the second order in $\partial_{k}Q$ and $\sigma$-matrices, under the constraint $Q^{2}=1$. Therefore, the structure of all terms in the action, including those attributed to SOC, can be reconstructed from the space-time symmetry arguments. However, such arguments still allow for arbitrary complex coefficients in front of the invariants, whereas the microscopically derived coefficients in Eq. (1) are quite specific – some of them are real, and some are purely imaginary. As we will see, this is controlled by another set of fundamental constraints unrelated to the space symmetry group of the system.
@@
Another condition restricting the form of the action follows from the fact that the expectation value of $Q$ equals to the contour quasiclassical Green’s function $g(t,t^{\prime})=\langle Q(t,t^{\prime})\rangle$. On the one hand, Hermitian conjugation of a chronologically ordered propagator reverses the time ordering to anti-chronological. On the other hand, in the Keldysh matrix representation, the transformation between chronological and anti-chronological contour Green’s functions can be realized as a matrix operation. This leads to the following fundamental symmetry condition of the action of NLSM (see SM Sec. B),
@@
$\mathcal{J}_{k}$ $=-Dg\nabla_{k}g+\frac{i}{16}\gamma_{kl}\varepsilon_{lij}\{[\sigma_{i},g],\sigma_{j}+g\sigma_{j}g\}\,,$ (8)
$\mathcal{T}$ $=-\frac{i}{8}\gamma_{kl}\varepsilon_{lij}[\{\partial_{k}g,g\sigma_{i}g\},\sigma_{j}]\,.$ (9)
@@
$\mathcal{J}_{k}$ $=-Dg_{0}\nabla_{k}\delta g+\gamma_{kj}\delta g^{j}_{t}\,,$ (10)
$\mathcal{T}$ $=\gamma_{kj}\sigma_{j}\partial_{k}\delta g_{s}\,.$ (11)
@@ Appendix A Charge conjugation symmetry
In the description of superconductivity, it is convenient to introduce the bispinor of Fermion fields $\Psi=(\psi_{\uparrow},\psi_{\downarrow},\psi^{\dagger}_{\downarrow},-\psi^{\dagger}_{\uparrow})^{T}$ [25, 29, 30]. The vector $\Psi$ contains $(\psi_{\uparrow},\psi_{\downarrow})$ and its time-reversed partner $(\psi_{\downarrow}^{\dagger},-\psi_{\uparrow}^{\dagger})$, where the time reversal operator is defined as $\sigma_{2}\mathcal{K}$. Here $\sigma_{2}$ is the second Pauli matrix in spin space. This is suitable for the description of superconductivity, which couples time-reversed pairs of electrons [53]. The theory is most conveniently described in terms of $\Psi$ and $\bar{\Psi}=\psi^{\dagger}\tau_{3}=(\psi_{\uparrow}^{\dagger},\psi_{\downarrow}^{\dagger},-\psi_{\downarrow},\psi_{\uparrow})$ instead of $\Psi$ and $\Psi^{\dagger}$, so that any term that is the same for electrons and holes comes with $\tau_{0}$, while any term that is opposite for electrons and holes comes with $\tau_{3}$.
Consequently, not all degrees of freedom are independent. Indeed, $\bar{\Psi}$ and $\Psi$ are related by $\Psi=C\bar{\Psi}^{T}$, and hence $\bar{\Psi}=\Psi C^{-1}$, where $C=-i\tau_{1}\sigma_{2}$ is the so-called charge conjugation symmetry operator. Therefore, the matrix composed of $\Psi\bar{\Psi}$ always satisfies the relation $C(\Psi(t^{\prime})\bar{\Psi}(t))^{T}C^{-1}=C\bar{\Psi}^{T}(t)\Psi^{T}(t^{\prime})C^{-1}=\Psi(t)\bar{\Psi}(t^{\prime})$. Notably, this holds for all combinations of Fermion fields $\Psi,\bar{\Psi}$ in the domain of integration.
Since the Q-matrix is used for Hubbard-Stratonovich decoupling of $\Psi\bar{\Psi}$ [29, 31], the domain of integration for the Q-matrix should be restricted to those Q-matrices that satisfy this same symmetry. In contrast to the main text, we will here distinguish the Q-matrices on the contour from those written as matrices in Keldysh space by using the notation $\mathcal{Q}$ on the contour and $Q$ in Keldysh space. On the contour this symmetry can be written in the same form as the symmetry on $\Psi\bar{\Psi}$:
@@
In the main text the non-linear sigma model is, as usual, presented in the rotated Keldysh space. First, the non-rotated Keldysh space is created by writing the different branches in matrix form, that is,
@@
where the superscripts $\pm$ have been used to distinguish between times taken on the forward and backward branches of the contour and $t,t^{\prime}$ are real times. The charge conjugation symmetry may be written in terms of this representation as
@@
The rotated Keldysh space can now be obtained via the matrix transformation
@@
where $L=1/\sqrt{2}(1-i\rho_{2})$ and $\rho_{i}$ is the $i$-th Pauli matrix in Keldysh space is performed because in this basis any causal function on the contour, such as the Green’s function, is upper triangular in Keldysh space [31, 54, 55, 30, 32]. In this rotated Keldysh space the constraint on the manifold of integration can be written as
@@
where the last equality follows from the fact that $C$ commutes with any Pauli matrix in $\rho$-space. We may now use $L\rho_{3}L^{-1}=\rho_{1}$ to find the charge conjugation relation on $Q$:
@@
This symmetry will be referred to as the charge conjugation symmetry.
@@ Appendix B Chronological symmetry
In this section we show that the condition $g=-\rho_{2}\tau_{3}g^{\dagger}\tau_{3}\rho_{2}$ and the consequent chronological symmetry of the action can be derived from the following two assumptions:
@@
Replacing $\tilde{Q}$ by $Q$ in the last expression we conclude that this we may impose the symmetry $iS[Q]=(iS[-\tau_{3}\rho_{2}Q^{\dagger}\rho_{2}\tau_{3}])^{*}$ on the action.
@@ Appendix C Terms combining Pauli matrices and derivatives
C.1 Second order: One Pauli matrix, one derivative
There exist two terms with one Pauli matrix and one derivative, $\sigma_{k}\partial_{j}Q$ and $\sigma_{k}Q\partial_{j}Q$. Applying transposition to the first we find
@@ C.1 Second order: One Pauli matrix, one derivative
Next, we consider the influence of the Hermitian conjugation symmetry. We find
@@
where $\tilde{Q}=-\rho_{2}\tau_{3}Q^{\dagger}\tau_{3}\rho_{2}$. Since $iS(Q)^{*}=iS(\tilde{Q})$ it now follows that $\tilde{\alpha}_{kj}=-\tilde{\alpha}_{kj}^{*}$, that is, the tensor has only imaginary indices. Thus, $\tilde{\alpha}_{kj}=i\alpha_{kj}$, where $\alpha_{kj}$ is a real second rank pseudotensor. The corresponding term in the bulk is
@@
In the Keldysh equation of the normal state this becomes $2i\alpha_{kj}[\sigma_{k},\partial_{j}F]$. Meanwhile its contribution to the current is
@@
C.2 Third order: Two Pauli matrices, one derivative
@@
that is, charge conjugation symmetry requires the corresponding tensor $\beta_{ijk}$ satisfies $\beta_{ijk}=\beta_{jik}$.
For chronological symmetry we calculate
@@
that is $\beta_{ijk}=\beta_{jik}^{*}=\beta_{ijk}$, where in the last equality we used charge conjugation symmetry. Thus, this tensor is real. The contribution to the bulk equation is
@@
$J_{k}^{K}=4\beta_{ijk}(h-\sigma_{j}h\sigma_{i})$ (38)
@@
that is, the third rank pseudotensor corresponding to this term is antisymmetric in the first two indices and we may write it as $\varepsilon_{ijl}\tilde{\gamma}_{lk}$, where $\tilde{\gamma}_{lk}$ is a second rank pseudotensor.
For chronological symmetry we take the Hermitian conjugation and calculate
@@
In the normal state this becomes
@@
C.3 Third order: One Pauli, two derivatives
@@
that is, $\tilde{\varkappa}_{ijk}=-\tilde{\varkappa}_{ikj}$ and we may write it as $i\frac{D}{4}\varepsilon_{jkl}\varkappa_{il}$. The factor $\frac{D}{4}$ was chosen so that the $\chi_{al}$ corresponds to the spin-swapping coefficient in isotropic materials, as shown in section C.3.1. chronological symmetry gives
@@
Thus, the pseudotensor of this term satisfies $\tilde{\theta}_{ijk}=-\tilde{\theta}_{ikj}$ and we may write the tensor belonging to this term as $-\frac{D}{4}\varepsilon_{jkl}\theta_{il}$. The choice for the factor $-\frac{D}{4}$ was made to ensure that that in an isotropic material our tensor $\theta_{il}$ reduces to the conventional spin-Hall angle, as shown in section C.3.1.
@@
C.3.1 Relation to usual definitions in isotropic materials
In isotropic materials, the only allowed third rank pseudotensor is the fully antisymmetric one, and hence we may write and $\theta_{il}=\theta\delta_{il}$ and $\varkappa_{il}=\varkappa\delta_{il}$. In this case the current in the normal state may be written as

### v2

@@ Abstract
Here we propose another approach that, similarly to the Ginzburg-Landau (GL) theory, relies solely on symmetry arguments. This allows us to incorporate effects of spin-orbit coupling into the NLSM and formulate a phenomenological quantum kinetic theory. Our theory captures all symmetry-allowed transport effects associated with spin, charge, and their coupling, such as spin-Hall and spin-galvanic effects. In contrast to GL theory, this phenomenology is valid for arbitrary temperatures, away from equilibrium, and covers the normal state. In the following, we identify the basic symmetries and explain how to construct the generalized NLSM. Then, the saddle point condition leads us to our main objective: obtaining the general quantum kinetic equation for disordered conductors with spin-orbit coupling, which reveals all physical effects associated with the different terms of the action.
@@
-- $\frac{D}{4}\varepsilon_{ijk}\big{[}\theta\sigma_{i}Q\partial_{j}Q\partial_{k}Q-i\varkappa\sigma_{i}\partial_{j}Q\partial_{k}Q\big{]}+\frac{\Gamma}{8}\sigma_{k}Q\sigma_{k}Q\Big{)},$ (1)
where the Pauli matrices $\sigma_{i}$ and $\tau_{i}$ span the spin and the Nambu spaces, respectively, $\hat{\omega}_{tt^{\prime}}=\partial_{t}\delta(t-t^{\prime})$ is the time derivative operator, $D$ is the diffusion coefficient, and $\hat{\Delta}=\tau_{2}|\Delta|e^{i\tau_{3}\varphi}$ is the superconducting order parameter matrix. The second line in Eq. (1) describes the effects of extrinsic SOC: the spin Hall effect, spin current swapping, and the spin relaxation, parameterized by the Hall angle $\theta$, the swapping coefficient $\varkappa$, and the spin relaxation rate $\Gamma$, respectively. The external electromagnetic potential is introduced through the replacement $\partial_{\mu}\mapsto\partial_{\mu}-i[A_{\mu}\tau_{3},.]$, and the Zeeman or exchange field is denoted by $\mathbf{h}$.
By inspecting the structure of $S[Q]$ in Eq. (1), we can observe the following. In general, the action represents a low-energy gradient expansion combined with an expansion in powers of spin Pauli matrices $\sigma_{i}$, and involves $\tau_{3}$ in the terms/fields breaking the time-reversal symmetry. Each term in $S$ is a scalar composed of $Q$, $\partial_{k}Q$, and $\sigma_{i}$, and respecting the spatial symmetry of the system. In fact, the terms present in Eq. (1) exhaust all scalar rotation invariant forms up to the second order in $\partial_{k}Q$ and $\sigma$-matrices, under the constraint $Q^{2}=1$. Therefore, the structure of all terms in the action, including those attributed to SOC, can be reconstructed from the space-time symmetry arguments. However, such arguments still allow for arbitrary complex coefficients in front of the invariants, whereas the microscopically derived coefficients in Eq. (1) are quite specific – some of them are real, and some are purely imaginary. As we will see, this is controlled by another set of fundamental constraints unrelated to the space symmetry group of the system.
@@
Another condition restricting the form of the action follows from the fact that the expectation value of $Q$ equals to the Keldysh contour quasiclassical Green’s function $g(t,t^{\prime})=\langle Q(t,t^{\prime})\rangle$. On the one hand, Hermitian conjugation of a chronologically ordered propagator reverses the time ordering to anti-chronological. On the other hand, in the Keldysh matrix representation, the transformation between chronological and anti-chronological contour Green’s functions can be realized as a matrix operation. This leads to the following fundamental symmetry condition of the action of NLSM (see SM Sec. B),
@@
$\mathcal{J}_{k}$ $=-Dg\nabla_{k}g+\frac{i}{16}\varepsilon_{ijl}\gamma_{lk}\{[\sigma_{i},g],\sigma_{j}+g\sigma_{j}g\}\,,$ (8)
$\mathcal{T}$ $=-\frac{i}{8}\varepsilon_{ijl}\gamma_{lk}[\{\partial_{k}g,g\sigma_{i}g\},\sigma_{j}]\,.$ (9)
@@
$\mathcal{J}_{k}$ $=-Dg_{0}\nabla_{k}\delta g+\gamma_{jk}\delta g^{j}_{t}\,,$ (10)
$\mathcal{T}$ $=\gamma_{jk}\sigma_{j}\partial_{k}\delta g_{s}\,.$ (11)
@@ Appendix A Charge conjugation symmetry
In the theoretical description of superconductivity, it is convenient to introduce the bispinor of Fermion fields $\Psi=(\psi_{\uparrow},\psi_{\downarrow},\psi^{\dagger}_{\downarrow},-\psi^{\dagger}_{\uparrow})^{T}$ [25, 29, 30]. The bispinor $\Psi$ contains both $(\psi_{\uparrow},\psi_{\downarrow})$ and its time-reversed partner $(\psi_{\downarrow}^{\dagger},-\psi_{\uparrow}^{\dagger})$, where the time reversal operator is defined as $\sigma_{2}\mathcal{K}$. Here $\sigma_{2}$ is the second Pauli matrix in spin space, and $\mathcal{K}$ denotes complex conjugation. This is suitable for the description of superconductivity, which couples time-reversed pairs of electrons [53]. The theory is most conveniently described in terms of $\Psi$ and $\bar{\Psi}=\psi^{\dagger}\tau_{3}=(\psi_{\uparrow}^{\dagger},\psi_{\downarrow}^{\dagger},-\psi_{\downarrow},\psi_{\uparrow})$ instead of $\Psi$ and $\Psi^{\dagger}$, so that any term that is the same for electrons and holes comes along with a $\tau_{0}$, while any term that is opposite for electrons and holes comes with $\tau_{3}$.
Consequently, not all degrees of freedom are independent. Indeed, $\bar{\Psi}$ and $\Psi$ are related by $\Psi=C\bar{\Psi}^{T}$, and hence $\bar{\Psi}=\Psi C^{-1}$, where $C=-i\tau_{1}\sigma_{2}$ is the so-called charge conjugation symmetry operator. Therefore, the matrix composed of $\Psi\bar{\Psi}$ necessarily satisfies the relation $C(\Psi(t^{\prime})\bar{\Psi}(t))^{T}C^{-1}=C\bar{\Psi}^{T}(t)\Psi^{T}(t^{\prime})C^{-1}=\Psi(t)\bar{\Psi}(t^{\prime})$. Notably, this holds for all combinations of Fermion fields $\Psi,\bar{\Psi}$ in the domain of integration.
Since the Q-matrix is used for Hubbard-Stratonovich decoupling of $\Psi\bar{\Psi}$ [29, 31], the domain of integration for the Q-matrix should be restricted to those Q-matrices that satisfy this same symmetry. In contrast to the main text, here we distinguish the Q-matrices on the contour from those written as matrices in Keldysh space by using the calligraphic $\mathcal{Q}$ on the contour and a non-calligraphic $Q$ in Keldysh space. On the contour the charge conjugation symmetry can be written in the same notation as the symmetry on $\Psi\bar{\Psi}$:
@@
In the main text the non-linear sigma model is, for convenience, presented in the rotated Keldysh space. First, the non-rotated Keldysh space is created by writing the different branches in matrix form, that is,
@@
where the superscripts $\pm$ have been used to distinguish between times taken on the forward and backward branches of the contour and $t,t^{\prime}$ are real times. The charge conjugation symmetry takes a form similar to that on the contour representation and may be written as
@@
The rotated Keldysh space can now be obtained via the transformation
@@
where $L=1/\sqrt{2}(1-i\rho_{2})$ and $\rho_{i}$ is the $i$-th Pauli matrix in Keldysh space is performed because in this basis any causal function on the contour, such as the Green’s function, is upper triangular in Keldysh space [31, 54, 55, 30, 32]. In this rotated Keldysh space the charge conjugation symmetry constraint on the manifold of integration can be written as
@@
where the last equality follows from the fact that $C$ commutes with any Pauli matrix in $\rho$-space. We may now use $L\rho_{3}L^{-1}=\rho_{1}$ to find the form charge conjugation relation on $Q$ presented in the main text:
@@
This symmetry is referred to as the charge conjugation symmetry and introduced in Eq. 2 in the main text.
@@ Appendix B Chronological symmetry
In this section we show that the condition $g=-\rho_{2}\tau_{3}g^{\dagger}\tau_{3}\rho_{2}$ and the consequent chronology symmetry of the action can be derived from the following two assumptions:
@@
Replacing $\tilde{Q}$ by $Q$ in the last expression we conclude that this we may impose the symmetry $iS[Q]=(iS[-\tau_{3}\rho_{2}Q^{\dagger}\rho_{2}\tau_{3}])^{*}$ on the action (cf. Eq. (3) in the main text).
@@ Appendix C Terms combining Pauli matrices and derivatives
In this section, we derive the different terms leading to Eqs. (4-6) in the main text. For this, we assume a weak SOC and determine the allowed terms in the action to second and third order in $\sigma$-matrices and derivatives.
C.1 Second order terms
Terms containing two Pauli matrices or two derivatives, provide the general relaxation, $\Gamma_{ij}$, and diffusion tensors, $D_{ij}$, in Eq. (4). Here we focus on the derivation of other terms that contain one Pauli matrix and one derivative. There are two of such terms, $\sigma_{k}\partial_{j}Q$ and $\sigma_{k}Q\partial_{j}Q$. Applying transposition to the first we find
@@ C.1 Second order terms
Next, we consider the influence of chronology symmetry. We find
@@
where $\tilde{Q}=-\rho_{2}\tau_{3}Q^{\dagger}\tau_{3}\rho_{2}$. Since $iS(Q)^{*}=iS(\tilde{Q})$ it now follows that $\tilde{\alpha}_{kj}=-\tilde{\alpha}_{kj}^{*}$, that is, the tensor has only imaginary indices. Thus, $\tilde{\alpha}_{kj}=i\alpha_{kj}$, where $\alpha_{kj}$ is a real second rank pseudotensor. The corresponding term in the bulk equation, which can be derived via the requirement that at the saddle point the action is invariant under variations $g+[\delta g,g]$, is
@@
In the Keldysh equation of the normal state this becomes $2i\alpha_{kj}[\sigma_{k},\partial_{j}F]$. Because the action contains terms with derivatives, integration by parts is necessary to derive the saddle point condition. This additionally yields terms that are evaluated only at the boundary. These terms have the form $n_{j}J_{j}$, where $n_{j}$ is the normal vector of the boundary and $J_{j}$ is the contribution to the current of the term under consideration. Equivalently, the expression for the current can be found by variation of the action with the vector potential. The contribution of the term $i\alpha_{kj}\sigma_{k}Q\partial_{j}Q$ to the current is
@@
C.2 Third order terms with two Pauli matrices and one derivative
@@
that is, charge conjugation symmetry requires that the corresponding tensor $\beta_{ijk}$ satisfies $\beta_{ijk}=\beta_{jik}$.
To find the restriction by chronological symmetry we calculate
@@
that is $\beta_{ijk}=\beta_{jik}^{*}=\beta_{ijk}$, where in the last equality we used the charge conjugation symmetry relation derived above. Thus, this tensor is real. The contribution to the bulk equation is
@@
$J_{k}^{K}=4\beta_{ijk}(F-\sigma_{j}F\sigma_{i})$ (38)
@@
that is, the third rank pseudotensor corresponding to this term is antisymmetric in the first two indices and we may write it as $\varepsilon_{ijl}\tilde{\gamma}_{lk}$, where $\varepsilon_{ijl}$ is the fully antisymmetric third rank pseudotensor and $\tilde{\gamma}_{lk}$ is a second rank pseudotensor.
For chronological symmetry we take the Hermitian conjugate and calculate
@@
In the normal state the Keldysh component of this expression becomes
@@
C.3 Third order terms with one Pauli matrix and two derivatives
@@
that is, $\tilde{\varkappa}_{ijk}=-\tilde{\varkappa}_{ikj}$ and we may write it as $i\frac{D}{4}\varepsilon_{jkl}\varkappa_{il}$, where $\varepsilon_{jkl}$ is the third rank antisymmetric pseudotensor and $\varkappa_{il}$ is a second rank tensor. The factor $\frac{D}{4}$ was chosen so that the $\chi_{al}$ corresponds to the spin-swapping coefficient in isotropic materials, as shown in section C.3.1. chronological symmetry gives
@@
Thus, the pseudotensor of this term satisfies $\tilde{\theta}_{ijk}=-\tilde{\theta}_{ikj}$ and we may write the tensor belonging to this term as $-\frac{D}{4}\varepsilon_{jkl}\theta_{il}$, where $\theta_{il}$ is a second rank tensor. The choice for the factor $-\frac{D}{4}$ was made to ensure that that in an isotropic material our tensor $\theta_{il}$ reduces to the conventional spin-Hall angle, as shown in section C.3.1.
@@
C.3.1 Isotropic materials
The above contributions to the matrix current, C22, C23, C27,C28, have been derived in the most general case. If the system is isotropic, the only allowed second rank tensor is the identity tensor, and hence we may write and $\theta_{il}=\theta\delta_{il}$ and $\varkappa_{il}=\varkappa\delta_{il}$. In this case the current in the normal state may be written as